ok please please help me on this one my brain is about to burst
Mr. Williams is three times as old as his son. Twelve years from now, he will be only twice as old as his son. What are their ages now? In 12 years?
Mr Williams now?
son now?
Mr. Williams in 12 years?
son in 12 years?

Found 2 solutions by Fermat, Paul:
Answer by Fermat(127)

let W represent the age of mr Williams
let S represent the age of mr Williams' son
Currently, mr williams is three times as old as his son: this becomes,
W = 3S
======
Twelve years from now, he will be only twice as old as his son. This becomes,
W + 12 = 2x(S + 12)
W + 12 = 2S + 24
W = 2S + 12
===========
We now substitute for W = 3S into the above eqn,
3S = 2S + 12
S = 12
======
W = 36
======

Answer by Paul(988)

Mr. Williams = 3x
His son = x

Twelve years from now his son's age = x+12
Twelve years from now his age = 3x+12

Equation:

2x+24=3x+12
-x=-12
x=12

3(12)=36

36+12=48
12+12=24

Hence, Mr. Williams is 36 years old and from twelve years from now he will be 48 years old and his son is 12 years old and from twelve years from now he will be 24 years old.
Paul.